Use the following amortization chart: Selling price of home Down payment Principal (loan) Rate of interest Years Payment per $1,000 Monthly mortgage payment $ 92,000 $ 6,000 $ 86,000 6% 30 $ 6.00 $ 516.00 Assume the interest rate rises to 7.5%. What is the total cost of interest with the new interest rate? (Use Table 15.1). (Do not round intermediate calculations. Round your final answer to the nearest cent.)
Answer #1

1. Formula for calculating monthly loan payment ( EMI )

EMI = { P × r × (1+r)n } ÷ { (1+r)n - 1 }

Here, EMI = monthly loan payment

P = loan amount = $86000

r = rate per period = 7.5 ÷ (12×100) = 0.00625

n = number of compounding periods = 30 × 12 = 360

Putting values in the formula we get

EMI = { 86000 × 0.00625 × (1+0.00625)360 } ÷ { (1+0.00625)360 - 1 } = $601.32

Total amount paid = 601.32 × 360 = $216475.20

Total interest paid = 216475.20 - 86000 = $130475.20